Mr. TP is a 68 year old of French descent. He spends quiet time with his dog in his little old house near the city. His wife passed away 10 years ago in a car accident. Fortunately, he has enough money to live well and stay in their house. Mr. TP’s two sons had to move out of town a year ago for work, but faithfully call him once a week. Mr. TP often feels lonely and isolated. For the last week his sons had not been able to reach him, even though they left numerous messages on his answering machine. They started to become worried after not talking to him for many days. They decided to come visit him and found him in bed. Mr. TP said he had been too tired to come out bed for the last 5 days, so they took him to the nearest hospital where he was admitted for an assessment.

His height and weight upon admission were 5’7” and 105 lbs.

  • Elbow breadth: 3.36 inches (74 mm)

Finally, you obtain the following information during an interview with Mr. TP:

He used to weight about 145 lbs., but he fell on an ice patch and broke his right leg last winter. Since then, he has had very little appetite, feels weak, and has lost weight progressively during the last 6 months. He stays around the house and calls the grocery store for home delivery once a week.

Mr. TP said that he had always been healthy before breaking his leg last winter. He never took any medications. He does not keep alcohol at home and he smokes a pipe of tobacco after dinner.

QUESTIONS:

  1. What is his ideal body weight, using the Hamwi method? Show your work.
  2. What is his weight change? Show your work.
  3. What is his percent weight change? Show your work and give your interpretation.
  4. What is his percent ideal body weight? Show your work and give your interpretation.
  5. Construct a PES statement for TP.
Answer #1

1. His ideal body weight using Hamwi method :

Hamwi method is based on inches and pounds, according to the following:

M = 106 lb for the first 5 ft + 6 lb for each additional inch.

M = 48 kg for the first 152.4 cm + 1.1 kg for each additional cm.

F = 100 lb for the first 5 ft + 5 lb for each additional inch.

F = 45 kg for the first 152.4 cm + 0.9 kg for each additional cm.

Weight of the patient is 105 lb

Height of the patient is 5 feet 7inches

So, his ideal weight according to Hamwi methlod is :

106 lb for first 5feet and 6lb additional for each inch = 6×7 lb = 106 +42 lb = 148lb is his ideal weight

2. The patient used to weight 145lb before fracture,but after fracture of his right leg his appetite decreased drastically and he now weights 105lb .His weight change is 145-105 = 40lbs he lost

3 His percentage of weight change is 40( weight loss ) / 145( his usual weight ) × 100

= 40/145× 100 = 27.5%

27.5% of his total weight (usual ) he has lost .